a triangular parking lot has two sides that are the same length and the third side is 5m longer. if the perimeter is 71m, find the lengths of the sides
Found 2 solutions by Adebimpe, shweta:
Answer by Adebimpe(3)

A triangle has 3 sides
the lenght of one side is 5cm
the other two sides are of the same lenght,: lets each be x, x respectively
the perimeter is 71cm
the perimeter is the addition of all the sides together; x + x + 5 = 71
2x = 71-5
2x = 66
x = 33cm

Answer by shweta(56)

Let the two equal sides of the triangular parking lot be x m
and the third side= x+5 m
Perimeter= 71m
Perimeter,P= x+ x+ x+5
71 = 3x+ 5
3x= 71-5
3x= 66
x= 66/3
x=22m
So two equal sides of the parking lot are 22m each and the third side is (22+5)=27 m